# Unlocking LLM Repair Capabilities in Low-Resource Programming Languages Through Cross-Language Translation and Multi-Agent Refinement

### v3

@@ 5.3.1. Experimental Design
On the other hand, Figure 6 presents the final $Pass@10$ improvements of the reasoning strategy on all languages. It is worth noting that Rust bugs benefit the most from translation-based repair, which achieves an increment of 22.09%. Following are the Go, PHP, Kotlin, C#, and Ruby bugs that undergo improvements of 11.31%, 9.33%, 8.87%, 8.82%, and 8.33%. Our statistical tests further confirm the improvements with a p-value of 0.0126 ($<0.05$) and a Cliff’s Delta of 0.64 (large effect).
@@
According to Figure 9(a), the reasoning strategy maintains higher $Precision@k$ until all strategies converge to the same level at the third iteration, demonstrating its early identification of valid fixes. Figure 9(b) shows that it achieves full recall with a $Recall@k$ of 1.00 by the fifth iteration, ahead of the greedy and random strategies, which converge only at the ninth and final iterations, respectively. Similarly, the $F1@k$ scores in Figure 9(c) highlight its early advantage, aligning with our previous $Pass@10$ observations.
[RQ-1] Findings: (1) The results of the greedy strategy suggest that simply defaulting to the language with the model’s historically best performance does not guarantee the most effective repairs across all cases. Instead, a more tailored approach that analyzes specific bugs to select the optimal target language leads to more robust repair performance. (2) LANTERN can significantly enhance program repair particularly for less common languages like Rust, achieving a maximal improvement of 22.09% in Pass@10. (3) The LLM’s reasoning capability can significantly enhance translation decision-making, with an average MAP@k of 0.957 and NDCG@k of 0.954, enabling efficient target language selection.
@@ 5.4.1. Experimental Design
Figure 11 exhibits the difficulty distribution of the bugs initially fixed by direct repair, fixed via translation, and those remain unfixed. We notice that the translation-based repair can fix bugs with 500 higher median difficulty for C#, while for Java, Kotlin, Ruby, and Rust, it enables successful repair of a group of bugs with median difficulty increased by 400. Our statistical analysis reveals a significant difference in difficulty between bugs fixed via translation and those initially fixed with a p-value of $485.25E-$ 48 and Cliff’s delta of 0.37 (medium effect).
@@
Figure 13. Examples of inconsistent transition.
For the inconsistent transitions, we identify that the inconsistency stems primarily from language-specific differences. Firstly, we notice that most bugs with COMPILATION ERROR can be fixed early in the translation phase, or transformed into other categories of bugs if minor additional issues remain within them. Figure 13 shows two examples of inconsistent transition. The first example presents a COMPILATION ERROR case of a C bug, which misses a semicolon ”;” is fixed after being translated to Python. Since Python syntax requires no semicolons, this compilation problem is naturally repaired after translation. We also find that such compilation errors can be fixed even when translating to languages that require semicolons, revealing the inherent features of LLM-based code translation. Secondly, language-specific differences such as memory and time usage are also critical contributors to the transition inconsistency. The second example in Figure 13 illustrates the transition of a MEMORY LIMIT EXCEEDED case of a C bug to a PASSED case in Rust, which consumes less memory resource, thus enabling the repair. We also notice that a very small subset of cases, where WRONG ANSWER cases are transformed into PASSED ones after translation also contribute to the inconsistency.
@@
We examine the consistency of back-translation in two granularity levels by comparing the bugs and samples before and after back-translation. Figure 14(a) shows the number of fixed bugs that remain correct and those that become incorrect after back-translation. We find that Java experienced the greatest loss with 11 out of 91 fixed bugs incorrectly translated, which remains a small number of bugs compared to its correctly translated bugs. Figure 14(b) exhibits the number of correct samples generated for different bugs before and after back-translation, among which C# suffers from a maximal loss of 182 correct samples, whereas Rust only encounters a minimal loss of six correct samples. Our statistical tests on correct samples before and after back-translation yield a p-value of 0.39 ($>0.05$) and a Cliff’s Delta of 0.22, confirming that the loss caused by back-translation is not significant.
[RQ-2] Findings: (1) The LLM-based reasoning can effectively identify optimal target languages where the model exhibits superior repair capabilities, achieving a minimal average translation path of 2.52. (2) The bugs addressed by LANTERN are complex and difficult to tackle, with some resolved bugs having median difficulty 400 to 500 larger. (3) Although language-specific errors sometimes introduce inconsistencies, translation generally preserves most bug semantics (with 12225 consistent PASSED cases and 3881 out of 4318 consistent WRONG ANSWER cases) and can even naturally repair certain defects. (4) Back-translation preserves semantic consistency for most repaired code, with an overall loss of only 7.56% of fixed bugs while 85.95% of correct samples are preserved.
@@ 5.5.2. Experimental Results
Figure 15 exhibits the $Pass@10$ results of the reasoning strategy and direct repair without translation. We notice that repair without translation shows only a slight performance improvement during the initial iteration, after which further fixes become consistently difficult to achieve. Notably, the reasoning strategy achieves a 16.55% improvement in $Pass@10$ for Rust compared to direct repair, followed by PHP, Go, and Ruby, with improvements of 7.27%, 6.96%, and 5.71%, respectively. Our statistical analysis of the final Pass@10 improvements confirms that translation-based repair significantly outperforms direct repair with a p-value of $031.95E-$ 03 and Cliff’s Delta of 0.42 (medium effect).
@@ 6. Threats to Validity
External. We evaluate on a large number of bugs (xCodeEval), demonstrating effectiveness across 11 diverse programming languages and bug types. Furthermore, our significant improvements on less common languages like Rust suggest that the approach generalizes beyond mainstream languages. Our approach can also potentially be applied to repository-level codebases. The pipeline’s design can also be extended to additional programming languages and models. To support open science, we have made our implementation publicly available.

### v4

@@ 5.3.1. Experimental Design
In RQ1.3, we compare LANTERN against SOTA approaches using DeepSeek-V3 on xCodeEval. Baselines include direct repair (original xCodeEval prompting), ChatRepair (iterative conversational refinement), Self-Planning (high-level algorithmic planning before coding), and Self-Collaboration (virtual development team with different agent roles for analysis, fixing, and validation).
@@
On the other hand, Figure 6 presents the final $Pass@10$ improvements of the reasoning strategy on all languages. It is worth noting that Rust bugs benefit the most from translation-based repair, which achieves an increment of 22.09%. Following are the Go, PHP, Kotlin, C#, and Ruby bugs that undergo improvements of 11.31%, 9.33%, 8.87%, 8.82%, and 8.33%. Our MWW tests (Mann and Whitney, 1947; Wilcoxon, 1992) further confirm the improvements with a p-value of 0.0126 (<0.05<0.05) and a Cliff’s Delta (Cliff, 1993; Yang et al., 2024a) of 0.64 (large effect).
@@
According to Figure 9(a), the reasoning strategy remains higher $Precision@k$ until all strategies converge to the same level at the third iteration, demonstrating its early identification of valid fixes. Figure 9(b) shows that it achieves full recall with a $Recall@k$ of 1.00 by the fifth iteration, ahead of greedy and random strategies, which converge only at the ninth and final iterations, respectively. Similarly, the $F1@k$ scores in Figure 9(c) highlight its early advantage, aligning with our previous $Pass@10$ observations.
5.3.4. Experimental Results for RQ1.3 (Approach Comparison)
Figure 10 shows LANTERN achieves the highest $Pass@10$ scores across xCodeEval, leading in 9 out of 11 languages. ChatRepair serves as a strong iterative baseline, demonstrating considerable gains from refinement within the original language (e.g., +18.31% for Rust). However, LANTERN consistently surpasses ChatRepair with improvements of 5.46%, 4.91%, and 6.76% on Go, Kotlin, and PHP respectively, suggesting that repetitive refinement has performance ceilings when confined to a single language. This gap stems from key architectural differences: ChatRepair iteratively refines solutions within one conversation, risking context accumulation and potential loss of critical early information, while LANTERN treats each repair as an independent session, ensuring context control and achieving one-shot solutions.
Self-Planning shows promise in some languages due to high-level planning abilities, achieving notable gains in JavaScript (+14.55%), but modest improvements elsewhere (e.g., +0.95% in C++) and significantly underperforming LANTERN. This indicates that generating high-level plans is insufficient when the LLM’s underlying implementation ability in specific languages is weak.
Self-Collaboration achieves large increments of 9.10% and 14.62% on JavaScript and Rust but causes severe performance degradation in several languages: C++ (-8.31%), Java (-12.13%), and most notably PHP (-26.48%). This suggests overly complex, simulated multi-role interactions can be inefficient, with the system potentially hallucinating incorrect analyses or validations that further mislead the model.
[RQ-1] Findings: (1) Simply defaulting to the language with the model’s historically best performance does not guarantee the most effective repairs. Instead, a more tailored approach that analyzes specific bugs to select the optimal target language leads to more robust repair performance. (2) LANTERN can significantly enhance program repair, particularly for less common languages like Rust, achieving a maximal improvement of 22.09% in Pass@10. (3) The LLM reasoning capability can significantly enhance translation decision-making, with an average MAP@k of 0.957 and NDCG@k of 0.954, enabling efficient target language selection. (4) Changing the linguistic context of a bug is a more powerful enhancement technique than simply iterating or reasoning more deeply within the original language, especially for complex bugs or in languages where the LLM’s native proficiency is limited.
@@ 5.4.1. Experimental Design
Figure 10. Pass@10 Performance of Different Approaches on xCodeEval
@@
Figure 12 exhibits the difficulty distribution of the bugs initially fixed by direct repair, fixed via translation, and those remain unfixed. We notice that the translation-based repair can fix bugs with 500 higher median difficulty for C#, while for Java, Kotlin, Ruby, and Rust, it enables successful repair of a group of bugs with median difficulty increased by 400. Our statistical analysis reveals a significant difference in difficulty between bugs fixed via translation and those initially fixed with a p-value of 5.25E485.25E-48 and Cliff’s delta of 0.37 (medium effect).
@@
First, we notice that most bugs with COMPILATION ERROR can be fixed early in the translation phase, or transformed into other categories of bugs if minor additional issues remain within them. Figure 14 shows two examples of inconsistent transition. The first example presents a COMPILATION ERROR case of a C bug, which misses a semicolon ”;” is fixed after being translated to Python. Since Python syntax requires no semicolons, this compilation problem is naturally repaired after translation. Such compilation errors can be fixed even when translating to languages that require semicolons, revealing the inherent features of LLM-based code translation. Secondly, language-specific differences such as memory and time usage are also critical contributors to the transition inconsistency. The second example in Figure 14 illustrates the transition of a MEMORY LIMIT EXCEEDED case of a C bug to a PASSED case in Rust, which consumes less memory resource, thus enabling the repair. A very small subset of cases, where WRONG ANSWER cases are transformed into PASSED ones after translation also contribute to the inconsistency.
Figure 15(a) shows the number of fixed bugs that remain correct and those that become incorrect after back-translation. Java experienced the greatest loss with 11 out of 91 fixed bugs incorrectly translated, which remains a small number of bugs compared to its correctly translated bugs. Figure 15(b) exhibits the number of correct samples generated for different bugs before and after back-translation, among which C# suffers from a maximal loss of 182 correct samples, whereas Rust only encounters a minimal loss of six correct samples. Our statistical tests on correct samples before and after back-translation yield a p-value of 0.39 (>0.05>0.05) and a Cliff’s Delta of 0.22, confirming that the loss caused by back-translation is not significant.
[RQ-2] Findings: (1) The LLM-based reasoning can effectively identify optimal target languages where model exhibits superior repair ability, achieving a minimal average translation path of 2.52. (2) The bugs addressed by LANTERN are complex and hard to tackle, with some resolved bugs having median difficulty 400 to 500 larger. (3) Although language-specific errors sometimes introduce inconsistencies, translation generally preserves most bug semantics (with 12225 consistent PASSED cases and 3881 out of 4318 consistent WRONG ANSWER cases) and can even naturally repair certain defects. (4) Back-translation preserves semantic consistency for most repaired code, with an overall loss of only 7.56% of fixed bugs while 85.95% of correct samples are preserved.
Table 2. Pass@10 Performance of Different LLMs on xCodeEval
LLM Approach C C# C++ Go Java JS Kotlin PHP Python Ruby Rust
Claude 3.5 Sonnet Direct 83.28 $82.68$ 75.30 82.29 79.81 79.80 75.94 $81.28$ 83.55 78.82 67.79
LANTERN 90.42 $87.34$ 80.81 90.46 85.99 86.40 92.89 $99.23$ 90.66 85.89 86.09
QWen2.5-72B-Instruct Direct 79.06 $69.50$ 65.90 74.97 67.19 82.61 84.47 $81.55$ 82.25 78.23 45.42
LANTERN 86.09 $79.92$ 75.26 85.11 76.89 88.30 88.21 $93.22$ 88.37 86.69 72.00
@@ 5.5.2. Experimental Results
Figure 16 exhibits the $Pass@10$ results of the reasoning strategy and direct repair without translation. We notice that repair without translation shows only a slight performance improvement during the initial iteration, after which further fixes become consistently difficult to achieve. Notably, the reasoning strategy achieves a 16.55% improvement in $Pass@10$ for Rust compared to direct repair, followed by PHP, Go, and Ruby, with improvements of 7.27%, 6.96%, and 5.71%, respectively. Our statistical analysis of the final Pass@10 improvements confirms that translation-based repair significantly outperforms direct repair with a p-value of 1.95E031.95E-03 and Cliff’s Delta of 0.42 (medium effect).
@@
5.6. RQ4: Generalizability
5.6.1. Experimental Design
In RQ4.1, we evaluate LANTERN on two real-world benchmarks. For Defects4J, we follow ChatRepair’s settings, focusing on the most comprehensive single-function subset (255 bugs from v1.2) and the filtered set of 78 bugs from v2.0. While ChatRepair used GPT-3.5, we employ DeepSeek-V3 for a more current comparison. For SWE-bench, we align with AGENTLESS evaluation settings, leveraging their fault localization results and extracting 10-line context windows around bug locations. We use AGENTLESS 1.5 with GPT-4o and employ the same model for fair comparison, measuring performance by resolved instances. Both benchmarks use the same 11 target languages from xCodeEval.
For RQ4.2, we investigate LANTERN’s generalizability across different models by replicating core xCodeEval experiments using Claude 3.5 Sonnet and Qwen2.5-72B-Instruct, with direct repair as the baseline. This determines whether our translation-based paradigm generalizes across model architectures.
5.6.2. Experimental Results for RQ4.1 (Real-world Generalizability)
Table 3. Performance on Defects4J and SWE-Bench Lite
Approach D4J1.2 D4J2.0 Approach SWE-Bench Lite
ChatRepair 141 54 AGENTLESSAGENTLESS 95/300 (31.67%)
LANTERN 170 60 LANTERNLANTERN 110/300 (36.67%)
Table 3 shows LANTERN’s performance on Defects4J and SWE-Bench compared to ChatRepair and AGENTLESS, respectively. On Defects4J 1.2, LANTERN resolves 170 bugs versus ChatRepair’s 141 (+29 bugs). For Defects4J 2.0, LANTERN fixes 60 bugs compared to ChatRepair’s 54, suggesting that while the source language is only Java, translating the buggy context to a different language (e.g., Python for its expressive syntax or C++ for its low-level control) appears to help the LLM overcome being stuck in a ”local optimum” of incorrect repair attempts and produce more effective patches.
On SWE-Bench Lite, LANTERN achieves stronger results, resolving 110/300 instances (36.67%) versus AGENTLESS’s 95/300 (31.67%). This benchmark requires resolving real-world GitHub issues across multiple files, yet LANTERN succeeds by translating only localized context windows around fault locations. This demonstrates that complete semantic representation of entire projects is unnecessary—focusing on relevant context maintains both scalability and efficiency for large-scale codebases.
The results indicate LANTERN’s effectiveness extends beyond traditionally under-represented languages. Even for widely-used languages like Java and Python, LANTERN transforms difficult instances into more tractable problems by leveraging different programming languages’ unique strengths, suggesting it functions as a general-purpose strategy for enhancing problem-solving capabilities on complex bugs.
5.6.3. Experimental Results for RQ4.2 (Model Generalizability)
Table 2 presents Pass@10 results for Claude 3.5 Sonnet and QWen2.5-72B-Instruct on xCodeEval. LANTERN elevates Claude’s performance across every language, with massive boosts in Kotlin (+16.94%), PHP (+17.95%), and Rust (+18.30%), demonstrating that even frontier proprietary models benefit significantly from cross-language translation.
Results with QWen2.5-72B-Instruct further highlight LANTERN’s power for open-source models. QWen’s baseline shows significant variance, particularly struggling with languages requiring strict compilation or complex type systems like Rust (45.42%) and C# (69.50%). LANTERN provides dramatic improvements on Rust (+26.58%), C# (+10.42%), Go (+10.13%), and PHP (+11.67%), revealing that LANTERN can elevate strong open-source models to achieve performance competitive with or superior to highly-capable proprietary models. For instance, QWen with LANTERN achieves 72.00% on Rust, surpassing Claude’s direct repair score of 67.79%.
[RQ-4] Findings: (1) LANTERN outperforms ChatRepair and AGENTLESS on the challenging real-world benchmarks, which demonstrates strong generalizabiltiy to real-world issue resolution, revealing its cross-language translation paradigm is effective for complex, real-world software defects, including those that require understanding context across multiple files. (2) LANTERN demonstrates model-agnostic generalizability across different LLMs. It consistently and significantly enhances the repair capabilities of both state-of-the-art closed-source and leading open-source LLMs.
@@ 6. Threats to Validity
External. We evaluate on a large number of bugs (xCodeEval), demonstrating effectiveness across 11 diverse programming languages and bug types. Furthermore, our significant improvements on less common languages like Rust suggest that the approach generalizes beyond mainstream languages. Our approach can also potentially be applied to repository-level codebases. The pipeline’s design can also be extended to additional programming languages and models. To support open science, we have made our implementation publicly available at https://github.com/stringing/LANTERN.